# Macroscopic degeneracy of ground state in the frustrated Heisenberg diamond chain

### v1

@@ II.1 General case
The excitation spectrum of the distorted diamond chain in the transition point is gapped. The gap in the one-magnon spectrum $\Delta E_{1}$ given by Eq.(9) for diamond in Fig.(3b) is $\Delta E_{1}=1/3$. Numerical calculations show that the minimal gap is $\Delta E\simeq 0.22$. Therefore, for $T<\Delta E$ the magnetic properties are determined by the contribution of the ground state manifold or, equivalently, it is given by that for the system of $n$ non-interacting spins $\frac{3}{2}$. In particular, the magnetization $m$ is $m=0$ at $T=0$ and it undergoes a jump from $m=\frac{1}{2}$ in the ferromagnetic phase to $m=0$ at the transition point. The susceptibility per spin $\chi$ diverges as $\chi\sim 1/T$ and the quantity $\chi T=\frac{5}{12}$ at $T=0$.
@@ II.3 Ideal diamond chain
For the ideal diamond chain the interactions $J_{i}$ are $J_{1}=J_{3}=J_{4}=J_{5}=-1$ and $J_{2}=1$ (see Fig.3c). In this special case there is a local conservation of the composite spin $\mathbf{L}_{i}=(\mathbf{\xi}_{i}+\mathbf{\eta}_{i})$ on the vertical diagonals of the diamonds. The composite spin is conserved quantity with quantum spin number $L_{i}=0$ or $L_{i}=1$. Then the model reduces to the spin chain with alternating spins $s=\frac{1}{2}$ and composite spins $L$, and it is described by the Hamiltonian
@@
All ground states of (36) can be enumerated as follows. First, we choose a definite configuration of $k$ singlets $L=0$ located in diamonds $\{i_{1},i_{2},\ldots i_{k}\}$. Number of such configurations with $k$ singlets is $C_{n}^{k}$. We note that singlet state on the diagonal of diamond is an exact state for the diamond, independent of the spin states on the left and the right sites of the diamond. Effectively, each singlet $L_{i}=0$ cuts the chain and creates open boundary at this place. Therefore, for chosen configuration of singlets $\{i_{1},i_{2},\ldots i_{k}\}$, we have $k$ sections of open chains, located between spins with $L=0$. The ground state of each of these open chains is the ferromagnetic state with all possible projections $S^{z}$, which can be written as
$\psi\left(S^{z}_{ij}\right)=\left|S_{max}(i,j),S^{z}_{ij}\right\rangle$ (37)
where the total spin of the ferromagnetic state of open diamond chain between the singlets on diagonals of $i$-th and $j$-th diamonds is $S_{max}(i,j)=\frac{3}{2}(j-i)-1$ and the spin projection can be $S^{z}_{ij}=-S_{max},\ldots S_{max}$. So that the number of multiplets of the ferromagnetic state between the singlets on $i$-th and $j$-th diamonds is $Q_{ij}=3(j-i)-1$.
@@
$\Psi\left(S^{z}_{i_{1}i_{2}},S^{z}_{i_{2}i_{3}},\ldots S^{z}_{i_{k}i_{1}}\right)=\left|s_{i_{1}}\right\rangle\psi\left(S^{z}_{i_{1}i_{2}}\right)\left|s_{i_{2}}\right\rangle\psi\left(S^{z}_{i_{2}i_{3}}\right)\ldots\left|s_{i_{k}}\right\rangle\psi\left(S^{z}_{i_{k}i_{1}}\right)$ (38)
where $\left|s_{i}\right\rangle$ means the singlet state on diagonal of $i$-th diamond. Therefore, the set of wave functions (38) contains $Q_{i_{1}i_{2}}\cdot Q_{i_{2}i_{3}}\cdot\ldots\cdot Q_{i_{k}i_{1}}$ states. For calculation of the total number of ground states one should sum up all such products for all possible singlet configurations $\{i_{1},i_{2},\ldots i_{k}\}$ with different $k$. The solution of this combinatorial problem is presented below.
@@
In this expression the first singlet is located in $k$-th diamond, therefore the remaining part contains $(n-k)$ diamonds with total number of states $w_{n-k}$ (see Fig.5). The factor $(3k-1)$ represents the number of multiplets of ferromagnetic state of $(k-1)$ diamonds, containing $(k-1)$ spins-$1$ and $k$ spins-$\frac{1}{2}$. The term with $k=n$ in the sum corresponds to the special case of just one singlet located on the last diamond. For this case there is just one spin-$\frac{1}{2}$ on the right side of the singlet and, therefore, we put $w_{0}=2$ for this case.
Figure 5: Scheme used for calculation of the ground state degeneracy for ideal diamond model.
The last term $(3n+2)$ in Eq.(39) corresponds to the case of no $L=0$ spins in the chain, and the degeneracy of this case is the number of multiplets of ferromagnetic state of the total open chain.
@@
For the calculation of the total number of degenerate states $W_{n}$ for the cyclic chain, containing $n$ diamonds we write equation
@@
The first term in Eq.(41) is a number of states when the last diamond (that loops the chain) contains diagonal singlet. The second term $(3n+1)$ in Eq.(41) corresponds to the case without singlets in the chain, and the degeneracy in this case is the number of multiplets of ferromagnetic state of the total system. The term $(n-1)(3n-1)$ corresponds to the case with one singlet on the whole chain. The sum contains terms when the last diamond belong to ferromagnetic section of length $m$.
@@
Similar approach can be used for the calculation of the ground state degeneracy in sectors with fixed total $S^{z}$. Let the total number of ground states of $n$ diamond in the sector with total $S^{z}=S_{\max}^{z}-m$ is $w_{n,m}$. Then $w_{n,m}$ can be expressed as a sum
@@
Here, we follow the method of calculating the ground state degeneracy used in Eq.(39) and shown in Fig.5. The solutions of Eq.(43) are
@@
Now let us consider two-magnon excitations. At first, we consider model (36) with only one spin $\mathbf{L}=0$ and all others $\mathbf{L}=1$. As mentioned earlier, the spin $\mathbf{L}=0$ cuts the chain so that initially cyclic chain effectively transforms to the open chain with $n$ spins $\frac{1}{2}$ (including ends) and $(n-1)$ spins $1$. Let us consider one-magnon excitations in this open chain, which correspond to the two-magnon excitations of initial cyclic chain. The one-magnon spectrum of open chain is given by Eq.(49), where $q$ is determined from equation
@@
The value of $q$ corresponding to the lowest excitation, $\Delta E_{2}$, is $q=\frac{\pi}{2(n-1)}$ at $n\gg 1$. Then $\Delta E_{2}=\frac{\pi^{2}}{6n^{2}}$, which is 4x lower than $\Delta E_{1}$. Numerical calculations show that $\Delta E_{2}$ is the lowest level in the entire excitation spectrum. Thus, the spectrum of the ideal diamond chain is gapless in the transition point and behaves as $\Delta E\sim n^{-2}$.
As well as for the distorted diamond chain in transition point, magnetization $m$ in the ideal diamond model is $m=0$ at $T=0$ and it undergoes a jump from $m=\frac{1}{2}$ in the ferromagnetic phase to $m=0$ at the transition point. The magnetization $m(h)$ at $T=0$ of the ideal diamond with $J_{2}=-J_{1}$ ($J_{1}<0)$ has been calculated in Gu for chain with 120 spins. The magnetization is a smooth curve from $m=0$ at $h=0$ to $m=\frac{1}{2}$ at the saturation magnetic field. The susceptibility diverges as $\chi\sim\frac{1}{T}$ and $\chi T=\frac{5}{12}$ at $T\to 0$ as for the distortion diamond chain. The specific heat $C(T)$ calculated in Gu has one maximum.
Deviation from the transition point leads to the ferromagnetic ground state for $J_{2}<1$ and for $J_{2}>1$ the ground state is $2^{n}$ degenerate and includes the states with $S_{tot}$ for $0\leq S_{tot}\leq\frac{n}{2}$. In terms of Hamiltonian (36) the ground state corresponds to all $L_{i}=0$ and $n$ free spins $s_{i}$. There is a finite gap to the state with $S_{tot}=\frac{n}{2}+1$, caused by the destruction of one of the singlets $L_{i}=0$.
@@ II.4 Diagonal state case
Let us consider the diamond model for which condition (10) is satisfied but (14) is not (a particular case is shown in Fig.3d). As was noted above, the exact one-magnon states are localized on vertical diagonals of the diamond chain and they are given by Eq.(11). The exact ground states in the spin sector $S_{tot}=S_{\max}-k$ $\ (0\leq k\leq n)$ can be written as a product of $k$ functions
@@
The number of the ground states $W(n,k)$ in the spin sector $S^{z}=S_{\max}-k$ is easily calculated
@@
The total number of the ground states $W(n)$ is
@@
The ground state degeneracy in this type of the diamond chain is exponentially large as well. It leads to the residual entropy $s_{0}=\frac{1}{3}\ln 2$ at $n\gg 1$. The magnetization curve has a form given by Eq.(18) and the magnetization at $h/T\to 0$ tends to $\frac{1}{3}$. Therefore, the diamond chain of this type is magnetically ordered at the transition point at $T=0$ and the magnetization similarly to the general case of the diamond chain undergoes a jump from $m=\frac{1}{2}$ in the ferromagnetic phase to $m=\frac{1}{3}$ at the transition point. As to the gap in the excitation spectrum it is $\sim\frac{1}{n^{2}}$ for one-magnon excitations (for example, $\Delta E_{1}$ in model with diamond in Fig.3d $\Delta E_{1}=\frac{5\pi^{2}}{6n^{2}})$ and generally goes down as $n^{-2}$ at $n\gg 1$ according to numerical calculations and, therefore, the spectrum is gapless.
@@ III Summary
We have studied the ground state of the spin-$1/2$ Heisenberg diamond chain with competing F and AF exchange interactions at the transition point between the ferromagnetic and other (singlet or ferrimagnetic) ground state phases. The ground state properties of this model in the transition point are highly nontrivial. The ground state consists of the localized magnons and special magnon complexes which form macroscopically degenerate ground state. We consider four types of the diamond chain. One of them is the diamond chain with only one-magnon localized states. The properties of this model are similar to that of the F-AF delta-chain. In particular, the ground state degeneracy $W(n)\simeq 2^{n}$ and the residual entropy is $\frac{1}{3}\ln 2$. The diamond chain of this type is magnetically ordered at $T=0$ and the magnetization undergoes a jump from $m=\frac{1}{2}$ in the ferromagnetic phase to $m=\frac{1}{3}$ at the transition point.
@@
The most striking feature of the second considered model, the distortion diamond chain, is the existence of two- and three-magnon localized states along with conventional one-magnon localized states. These states form flat bands in two- and three magnon spectrum and they are ground states as well. It turns out that the number of the ground state in the sectors with fixed total spin projection $S^{z}$ coincides with that for the system of $n$ non-interacting spins $\frac{3}{2}$. We have checked this fact by numerical calculations of finite chains. The localized magnon states in the distortion diamond chain lead to the ground state degeneracy $W(n)=4^{n}$, which exceeds exponentially that in the model with only one-magnon localized states. As a result, the residual entropy per spin is $s_{0}=\frac{2}{3}\ln 2$. The ideal diamond chain has the same residual entropy. For the ideal diamond chain, analytical results have been obtained for the number of the ground states in different sectors of $S^{z}$ and they are confirmed by numerical calculations of finite chains. The spectrum of excitations in three variants of diamond chains is gapless and it is gapped in the distorted diamond chain. The magnetization $m=0$ at $T=0$ in the transition point for the distorted and ideal diamond chains and, therefore, there is a magnetization jump from $m=\frac{1}{2}$ in the ferromagnetic phase to $m=0$ at the transition point. The susceptibility per spin $\chi$ diverges as $\chi\sim 1/T$ and the product $\chi T=\frac{5}{12}$ at $T=0$.
@@
In this paper we consider the isotropic diamond chain on the boundary of the ferromagnetic ground state. This model can be extended to the case of anisotropic interactions similar to that developed for F-AF delta chain in Ref.anis (details will be presented elsewhere).
@@
It is known that the spin systems with macroscopic degenerate ground state show enhanced magnetocaloric effect. From this point of view the considered diamond chain can serve as a beacon for the search for materials for low-temperature cooling. Of course, macroscopic degeneracy occurs at certain relations between exchange interactions, which can hardly be exactly satisfied in real compounds. However, some important features of the model remain if the values of interactions are close to the critical ones.
The numerical calculations were carried out with use of the ALPS libraries alps .

### v2

@@ II.1 General case
Figure 8: Dependence of the energy gap on the system size for distorted diamond case shown in Fig.3b, calculated numerically for $N=12,15,18,24$.
The excitation spectrum of the distorted diamond chain at the transition point is gapped. The gap in the one-magnon spectrum $\Delta E_{1}$, given by Eq.(9) for the diamond in Fig.3b, is $\Delta E_{1}=1/3$. Numerical calculations show that the minimal gap is lower, but still finite $\Delta E\simeq 0.22$ (see Fig.8). Therefore, for $T<\Delta E$ the system properties are determined by the contribution of the ground state manifold or, equivalently, it is given by the one for the system of $n$ non-interacting spins $\frac{3}{2}$.
Figure 9: Temperature dependencies of the specific heat (solid line), entropy (dashed line) and the product of magnetic susceptibility and temperature, $\chi T$, (dotted line) for the distorted diamond case shown in Fig.3b, calculated numerically for $N=18$. For better visibility the specific heat is multiplied by a factor of 3, $3C(T)$.
The thermodynamic properties of the distorted diamond chain shown in Fig.3b are demonstrated in Fig.9. As it is seen in Fig.9, the entropy is equal to its residual value (37) at low temperature $T<\Delta E$ and tends to its high temperature limiting value $\ln 2$. The specific heat $C(T)$ has one maximum with a wide left shoulder. The susceptibility per spin for $T<\Delta E$ behaves as for the system of non-interacting spins $\frac{3}{2}$: $\chi=\frac{5}{12T}$. For high temperature the susceptibility tends to $\chi=\frac{1}{4T}$.
Figure 10: Magnetization curve as a function of $h/T$ for distorted diamond case shown in Fig.3b, calculated numerically for $N=18$ and $T=0.1$ (dashed line) and $T=1$ (dotted line). Magnetization curve (38) is shown by solid line.
The magnetization per spin of the system of non-interacting spins $\frac{3}{2}$ is given by
$m=\frac{3\sinh(\frac{3h}{2T})+\sinh(\frac{h}{2T})}{6\cosh(\frac{3h}{2T})+6\cosh(\frac{h}{2T})}$ (38)
and is shown in Fig.10 by solid line. As one can see in Fig.10, the magnetization curve for $T=0.1$, which is lower than the energy gap, perfectly coincides with Eq.(38). But the magnetization curve for relatively high temperature $T=1$ substantially deviates from Eq.(38).
@@ II.3 Ideal diamond chain
For the ideal diamond chain the interactions $J_{i}$ are $J_{1}=J_{3}=J_{4}=J_{5}=-1$ and $J_{2}=1$ (see Fig.3c). In this special case there is a local conservation of the composite spin $\mathbf{L}_{i}=(\mathbf{\xi}_{i}+\mathbf{\eta}_{i})$ on the vertical diagonals of the diamonds. The composite spin is a conserved quantity with quantum spin number $L_{i}=0$ or $L_{i}=1$. Then the model reduces to the spin chain with alternating spins $s=\frac{1}{2}$ and composite spins $L$, and it is described by the Hamiltonian
@@
All ground states of (39) can be enumerated as follows. First, we choose a definite configuration of $k$ singlets $L=0$ located in diamonds $\{i_{1},i_{2},\ldots i_{k}\}$. The number of such configurations with $k$ singlets is $C_{n}^{k}$. We note that the singlet state on the diagonal of the diamond is an exact state for this diamond, independent of the spin states on the left and the right sites of it. Effectively, each singlet $L_{i}=0$ cuts the chain and creates an open boundary at this place. Therefore, for the chosen configuration of singlets $\{i_{1},i_{2},\ldots i_{k}\}$, we have $k$ sections of open chains, located between spins with $L=0$. The ground state of each of these open chains is the ferromagnetic state with all possible projections $S^{z}$, which can be written as
$\psi\left(S_{ij}^{z}\right)=\left|S_{max}(i,j),S_{ij}^{z}\right\rangle$ (40)
where the total spin of the ferromagnetic state of the open diamond chain between the singlets on diagonals of $i$-th and $j$-th diamonds is $S_{max}(i,j)=\frac{3}{2}(j-i)-1$ and the spin projection can be $S_{ij}^{z}=-S_{max},\ldots S_{max}$. Consequently, the number of multiplets of the ferromagnetic state between the singlets on $i$-th and $j$-th diamonds is $Q_{ij}=3(j-i)-1$.
@@
$\Psi\left(S_{i_{1}i_{2}}^{z},S_{i_{2}i_{3}}^{z},\ldots S_{i_{k}i_{1}}^{z}\right)=\left|s_{i_{1}}\right\rangle\psi\left(S_{i_{1}i_{2}}^{z}\right)\left|s_{i_{2}}\right\rangle\psi\left(S_{i_{2}i_{3}}^{z}\right)\ldots\left|s_{i_{k}}\right\rangle\psi\left(S_{i_{k}i_{1}}^{z}\right)$ (41)
where $\left|s_{i}\right\rangle$ means the singlet state on diagonal of $i$-th diamond. Therefore, the set of wave functions (41) contains $Q_{i_{1}i_{2}}\cdot Q_{i_{2}i_{3}}\cdot\ldots\cdot Q_{i_{k}i_{1}}$ states. To calculate the total number of ground states one should sum up all such products for all possible singlet configurations $\{i_{1},i_{2},\ldots i_{k}\}$ with different $k$. The solution of this combinatorial problem is presented below.
@@
In this expression the first singlet is located in $k$-th diamond, therefore the remaining part contains $(n-k)$ diamonds with the total number of states $w_{n-k}$ (see Fig.11). The factor $(3k-1)$ represents the number of multiplets of the ferromagnetic state of $(k-1)$ diamonds, containing $(k-1)$ spins-$1$ and $k$ spins-$\frac{1}{2}$. The term with $k=n$ in the sum corresponds to the special case of just one singlet located on the last diamond. For this case there is just one spin-$\frac{1}{2}$ on the right side of the singlet and, therefore, we put $w_{0}=2$ for this case.
Figure 11: Scheme used for the calculation of the ground state degeneracy for an ideal diamond model.
The last term $(3n+2)$ in Eq.(42) corresponds to the case of no $L=0$ spins in the chain, and the degeneracy of this case is the number of multiplets of the ferromagnetic state of the total open chain.
@@
For the calculation of the total number of degenerate states $W_{n}$ for the cyclic chain, containing $n$ diamonds we write the equation
@@
The first term in Eq.(44) is the number of states when the last diamond (that loops the chain) contains the diagonal singlet. The second term $(3n+1)$ in Eq.(44) corresponds to the case without singlets in the chain, and the degeneracy in this case is the number of multiplets of the ferromagnetic state of the total system. The term $(n-1)(3n-1)$ corresponds to the case with one singlet on the whole chain. The sum contains terms when the last diamond belongs to the ferromagnetic section of length $m$.
@@
A similar approach can be used for the calculation of the ground state degeneracy in sectors with fixed total $S^{z}$. Let the total number of ground states of $n$ diamonds in the sector with total $S^{z}=S_{\max}^{z}-m$ be $w_{n,m}$. Then $w_{n,m}$ can be expressed as a sum
@@
Here, we follow the method of the calculation of the ground state degeneracy used in Eq.(42) and shown in Fig.11. The solutions of Eq.(46) are
@@
Now let us consider two-magnon excitations. At first, we consider the model (39) with only one spin $\mathbf{L}=0$ and all others $\mathbf{L}=1$. As mentioned earlier, the spin $\mathbf{L}=0$ cuts the chain so that the initially cyclic chain effectively transforms to the open chain with $n$ spins $\frac{1}{2}$ (including ends) and $(n-1)$ spins $1$. Let us consider one-magnon excitations in this open chain, which correspond to the two-magnon excitations of initial cyclic chain. The one-magnon spectrum of the open chain is given by Eq.(52), where $q$ is determined from the equation
@@
The value of $q$ that corresponds to the lowest excitation, $\Delta E_{2}$, is $q=\frac{\pi}{2(n-1)}$ at $n\gg 1$. Then $\Delta E_{2}=\frac{\pi^{2}}{6n^{2}}$, which is 4x lower than $\Delta E_{1}$. Numerical calculations show that $\Delta E_{2}$ is the lowest level in the entire excitation spectrum. Thus, the spectrum of the ideal diamond chain is gapless in the transition point and behaves as $\Delta E\sim n^{-2}$.
As well as for the distorted diamond chain at the transition point, magnetization in the ideal diamond model is zero at $T=0$ and it undergoes a jump from $m=\frac{1}{2}$ in the ferromagnetic phase to $m=0$ at the transition point. The susceptibility and the specific heat of the ideal diamond model has been calculated in [27] for a chain with 120 spins. It was shown that the susceptibility diverges as $\chi\sim\frac{1}{T}$ and $\chi T=\frac{5}{12}$ at $T\to 0$ as for the distorted diamond chain. The specific heat $C(T)$ calculated in [27] has one maximum.
Deviation from the transition point leads to the ferromagnetic ground state for $J_{2}<1$ and for $J_{2}>1$ the ground state is $2^{n}$ degenerate and includes the states with $S_{tot}$ for $0\leq S_{tot}\leq\frac{n}{2}$. In terms of the Hamiltonian (39) the ground state corresponds to all $L_{i}=0$ and $n$ free spins $s_{i}$. There is a finite gap to the state with $S_{tot}=\frac{n}{2}+1$, caused by the destruction of one of the singlets $L_{i}=0$.
@@ II.4 Diagonal state case
Let us consider the diamond model for which the the condition (10) is satisfied but (14) is not (a particular case is shown in Fig.3d). As it was noted above, the exact one-magnon states are localized on vertical diagonals of the diamond chain and they are given by Eq.(11). The exact ground states in the spin sector $S_{tot}=S_{\max}-k$ $\ (0\leq k\leq n)$ can be written as a product of $k$ functions
@@
The number of ground states $W(n,k)$ in the spin sector $S^{z}=S_{\max}-k$ is equal to
@@
The total number of ground states $W(n)$ is
@@
The ground state degeneracy in this type of the diamond chain is exponentially large as well. In accordance with Eq.(18), this leads to the residual entropy $s_{0}=\frac{1}{3}\ln 2$ at $n\gg 1$.
Figure 12: Temperature dependencies of the specific heat (solid line), entropy (dashed line) and the product of magnetic susceptibility and temperature, $\chi T$, (dotted line) for the diagonal state case shown in Fig.3d, calculated numerically for $N=18$. For better visibility the specific heat is multiplied by a factor of 3, $3C(T)$.
The thermodynamic properties of the diagonal state chain shown in Fig.3d are demonstrated in Fig.12. As it is seen in Fig.12, the entropy tends to finite value defined by Eqs.(18),(57) at low temperatures $T<0.03$, and approaches to $\ln 2$ at high temperatures. The specific heat $C(T)$ has two maxima. The magnetic properties of the diagonal state chain is similar to the general case. The susceptibility diverges at $T\to 0$ faster than $1/T$. The magnetization curve in the thermodynamic limit has a form given by Eq.(19). Therefore, the diamond chain of this type is magnetically ordered at the transition point at $T=0$ and the magnetization in similarity to the general case of the diamond chain undergoes a jump from $m=\frac{1}{2}$ in the ferromagnetic phase to $m=\frac{1}{3}$ at the transition point.
Figure 13: Dependence of the energy gap on the system size for diagonal state case shown in Fig.3d, calculated numerically for $N=12,18,24,30$ (exact diagonalization and Lanczos algorithm). Dotted line is a fitting curve $\Delta E=\frac{3.12}{n^{2}}-\frac{3.8}{n^{3}}$.
The gap in one-magnon excitation spectrum (9) is equal to $\Delta E_{1}=\frac{5\pi^{2}}{6n^{2}}$ for the chain consisting of diamonds shown in Fig.3d. This means that the excitation gap goes down as $n^{-2}$ or faster at $n\gg 1$. The numerical calculations of the gap shown in Fig.13 confirm this law, and give $\Delta E=\frac{3.12}{n^{2}}$.
@@ III Summary
We have studied the ground state of the spin-$1/2$ Heisenberg diamond chain with competing F and AF exchange interactions at the transition point between the ferromagnetic and other (singlet or ferrimagnetic) ground state phases. The ground state properties of this model at the transition point are highly nontrivial. The ground state consists of the localized magnons and special magnon complexes which form macroscopically degenerate ground state. We consider four types of diamond chains. One of them is the diamond chain with only one-magnon localized states. The properties of this model are similar to those of the F-AF delta-chain. In particular, the ground state degeneracy is $W(n)\simeq 2^{n}$ and the residual entropy is $\frac{1}{3}\ln 2$. The diamond chain of this type is magnetically ordered at $T=0$ and the magnetization undergoes a jump from $m=\frac{1}{2}$ in the ferromagnetic phase to $m=\frac{1}{3}$ at the transition point.
@@
The most striking feature of the second considered model, the distorted diamond chain, is the existence of two- and three-magnon localized states along with the conventional one-magnon localized states. These states form flat bands in two- and three magnon spectrum and they are ground states as well. It turns out that the number of ground states in the sectors with fixed total spin projection $S^{z}$ coincides with the one for the system of $n$ non-interacting spins $\frac{3}{2}$. We have checked this fact by numerical calculations of finite chains. The localized magnon states in the distorted diamond chain lead to the ground state degeneracy $W(n)=4^{n}$, which exceeds exponentially the one in the model with only one-magnon localized states. As a result, the residual entropy per spin is $s_{0}=\frac{2}{3}\ln 2$. The ideal diamond chain has the same residual entropy. For the ideal diamond chain, analytical results have been obtained for the number of the ground states in different sectors of $S^{z}$ and they are confirmed by numerical calculations of finite chains. The spectrum of excitations in the three types of diamond chains is gapless and it is gapped in the distorted diamond chain. The magnetization is zero at $T=0$ at the transition point for the distorted and ideal diamond chains and, therefore, there is a magnetization jump from $m=\frac{1}{2}$ in the ferromagnetic phase to $m=0$ at the transition point. The susceptibility per spin $\chi$ diverges as $\chi\sim 1/T$ and the product $\chi T$ is $\chi T=\frac{5}{12}$ at $T=0$. The main features of four cases of the diamond model are summarized in the Table.1.
Table 1: Properties of four cases of diamond model.
General Distorted Ideal Diagonal
Ground state degeneracy $2^{n}+(n+4)C_{n}^{n/2}$ $4^{n}$ $14^{n}+3n-$ 1 $(2n+1)2^{n}$
Residual entropy per spin $\frac{1}{3}\ln 2$ $\frac{2}{3}\ln 2$ $\frac{2}{3}\ln 2$ $\frac{1}{3}\ln 2$
Lowest excitation, $\Delta E$ gapless, $\sim e^{-cn}$ gapped gapless, $\sim n^{-2}$ gapless, $\sim n^{-2}$
Low-T susceptibility, $\chi T$ diverges $\frac{5}{12}$ $\frac{5}{12}$ diverges
Specific heat two maxima one maximum one maximum two maxima
@@
In this paper we consider the isotropic diamond chain on the boundary of the ferromagnetic ground state. This model can be extended to the case of anisotropic interactions similar to those developed for F-AF delta chain in Ref.[32, *anis2, *anis3] (details will be presented elsewhere).
@@
It is known that spin systems with the macroscopic degenerate ground state show enhanced magnetocaloric effect. From this point of view the considered diamond chain can serve as a beacon for the search for materials for low-temperature cooling. Of course, macroscopic degeneracy occurs for certain relations between exchange interactions, which can hardly be satisfied exactly in real compounds. However, some important features of the model remain if the values of interactions are close to the critical ones. Recent progress in the synthesis of diamond chain compounds allows us to expect that compounds with parameters close to those given in this work can be synthesized in the near future.
Acknowledgements.
The numerical calculations were carried out with use of the ALPS libraries [35]. This work was financially supported by the Ministry of Sciences and Higher Education, Russian Federation (Research theme state registration number 122041400110-4).